Grade 12Physical Chemistry

equal moles of benzene and toulene are mixed the VP of benzene and toulene in pure state are 700 and 600 mm Hg respecti.The mole fraction of bnzene in vapour state is.?

Shriya Mehrotra
11 Years agoGrade 12

1 Answer

parthasarathy
11 Years ago
mole fraction=700/(700+600) =7/13.
but i am not sure